OpenStudy (anonymous):

write a quadratic equation having the given numbers as solutions -5 and -2

jimthompson5910 (jim_thompson5910):

x = -5 or x = -2 x+5 = 0 or x + 2 = 0 (x+5)(x+2) = 0 x^2 + 7x + 10 = 0

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (anonymous):

(x+5)(x+2) x^2 + 7x + 10 = 0